### v2

# CiteBART: Learning to Generate Citations for Local Citation Recommendation

Abstract
Local citation recommendation (LCR) suggests a set of papers for a citation placeholder within a given context. The task has evolved as generative approaches have become more promising than the traditional pre-fetch and re-rank-based state-of-the-art approaches. This paper introduces citation-specific pre-training within an encoder-decoder architecture, where author-date citation tokens are masked to learn to reconstruct them to fulfill LCR. There are two variants for this pre-training. In the local context-only base scheme (CiteBART-Base), the citation token in a local context is masked to learn to predict the citation. The global version (CiteBART-Global) extends the local context with the citing paper’s title and abstract to enrich the learning signal. CiteBART-Global achieves state-of-the-art performance on LCR benchmarks except for the FullTextPeerRead dataset, which is quite small to see the advantage of generative pre-training. The effect is significant in the larger benchmarks, e.g., Refseer and ArXiv., with the Refseer benchmark-trained model emerging as the best-performing model. We perform comprehensive experiments, including an ablation study, a qualitative analysis, and a taxonomy of hallucinations with detailed statistics. Our analyses confirm that CiteBART-Global has a cross-dataset generalization capability; the macro hallucination rate (MaHR) at the top-3 predictions is 4%, and when the ground-truth is in the top-k prediction list, the hallucination tendency in the other predictions drops significantly. We publicly share our code [^1] https://github.com/eyclk/CitationRecommendation, base datasets [^2] https://drive.google.com/drive/folders/1WlqlTkSj8LwihbrQvBX5F9_0uZAGGhiE?usp=drive_link, global datasets [^3] https://drive.google.com/drive/folders/1JH34nEXt8_p-0P9A--aQHK4yBXQfJe4v?usp=drive_link, and pre-trained models [^4] https://drive.google.com/drive/u/2/folders/1OBg6W3kQw4VWPMfrXEPxN8LzTopR1jak to support reproducibility.
Keywords Citation masking  $\cdot$ BART pre-training  $\cdot$ Local citation recommendation
1 Introduction
Citations are essential building blocks in scientific writing. Their accurate placements indicate quality, as one should know the literature to claim contributions and put the current study in the context of the existing work from different aspects, such as background information, method, and result comparison (Cohan et al., 2019).
Citation prediction is defined as a two-step process where the former focuses on where in the sentence to place the citation (Buscaldi et al., 2024), while the latter (citation recommendation) obtains a set of candidate papers once there is a specified citation placeholder in a given context. In this sense, citation recommendation serves as a citation suggestion mechanism. For a given scientific text, it can suggest additional papers on a similar topic. These suggestions can be considered additional reading material alongside the targeted paper, corresponding to the ground-truth citation.
There are two levels of citation recommendation: the first, whom to cite, and the second, whom to cite in what context. The former is global citation recommendation, traditionally performed based on paper metadata such as author names, paper titles, abstracts, conference venues, publisher information, etc. Recently, custom citation-aware language models (SciBERT (Beltagy et al., 2019), SPECTER (Cohan et al., 2020)) learn good citation-aware embeddings for full papers to perform well in this task. The latter task is local citation recommendation (LCR), aiming to determine the target paper for a citation placeholder.
LCR has been addressed in a few works. BERT-GCN (Jeong et al., 2020) utilizes a feedforward neural network to combine local citation context representations using BERT with citation encodings through Graph Convolutional Neural Networks (GCN). The most recent solutions to the problem adopt a two-step process that consists of pre-fetching and re-ranking. DualEnh (Medić and Snajder, 2020) enhances a local citation context with the citing article’s title and abstract and uses this enhanced context as the query vector to retrieve the most similar candidate articles using their titles and abstracts. It performs this ranking through BiLSTM representations of inputs with attention layers on top. On the other hand, HAtten (Gu et al., 2022) initially pre-fetches a set of papers using the nearest neighbor search between local citation context extended with the citing paper’s title and abstract (query text as a whole) and the title and abstracts from a given pool of papers. Afterward, it re-ranks the selected candidate papers using a fine-tuned SciBERT (Beltagy et al., 2019) model where the input is the query text concatenated with a candidate paper’s title and abstract.
The existing LCR works are not built upon Transformers but benefit from it indirectly, such as re-ranking the results (using fine-tuned SciBERT). Distinctively, we propose CiteBART, a custom pre-training approach based on the Transformer architecture. We mask citation tokens in the local contexts to learn to reconstruct them effectively during pre-training.
Fierro et al. (2024) support information-seeking using query-focused summarization, responding to user queries by answers with source attributions. Attributions are in the form of in-line references to the passages. In a similar direction, the ALCE benchmark (Gao et al., 2023) collects a diverse set of questions and retrieved passages to support answer generation with appropriate citations. As these models exhibit citation generation abilities, they are similar to CiteBART. However, CiteBART aims to fill in a passage with citations instead of targeting retrieval-based summaries with citations. In ALCE, there is a closed-book configuration where the model does not access any retrieved document to generate answers to a user query but is still different as the focus is answering a query instead of filling in the citation placeholder.
Table 1: An example for input and target formats for pre-training and evaluation with CiteBART. Due to space constraints, we present the contexts and abstracts in an abbreviated form.
Strategy Input Target
Base … error rate of 5.8% and a word error rate of 28.7%, which are on par with previous reported results <mask> . Unlike prior work, we do not use a language model during decoding and … Yao and Zweig, 2015
Global … error rate of 5.8% and a word error rate of 28.7%, which are on par with previous reported results <mask> . Unlike prior work, we do not use a language model during decoding and … </s> Deep Voice: Real-time Neural Text-to-Speech </s> We present Deep Voice, a production-quality text-to-speech system constructed entirely from deep neural … Yao and Zweig, 2015
In our approach, the base scheme (CiteBART-Base) learns through the masked citation context. In a second technique (CiteBART-Global), we extend the masked context with the citing paper’s global information, e.g., title and abstract (Table 1). Inspiring from pre-training under the REALM framework (Guu et al., 2020), we append this global information to the local context, allowing backpropagation through the global information to learn associations with the pool of papers from the corpus.
CiteBART achieves superior performance without relying on a pre-fetch and re-rank pipeline. It is an end-to-end learning system. On the other hand, a pre-fetch and re-rank pipeline, such as HAtten, utilizes the citing papers’ titles and abstracts with the local contexts to form the query encoding and the titles and abstracts of cited papers for the candidate papers’ representations. Thus, it exploits the titles and abstracts of papers from the test set to determine the cited papers for a citation placeholder. On the contrary, we do not use the global information (titles and abstracts) of target papers to make the recommendation. CiteBART-Global learns solely from the relation of citing papers’ global information with local citation contexts. The underlying assumption is one can find out the cited papers from the enhanced citation contexts, citation contexts that are concatenated with citing papers’ titles and abstracts. In the test phase, we feed these enhanced contexts to predict the target papers to be cited.
We summarize our contributions as follows:
  • We propose an end-to-end learning system, CiteBART, with custom citation masking for LCR.
  • CiteBART-Global achieves state-of-the-art performance on LCR benchmarks except for the FullTextPeerRead dataset, which is quite small to see the advantage of generative pre-training. The effect is significant in the larger benchmarks, e.g., Refseer and ArXiv. CiteBART-Base is still a strong baseline.
  • We provide a qualitative analysis to gain insight into the working of the approach, including the cross-dataset generalization capability.
  • We provide a taxonomy of hallucinated citations and report macro hallucination rates (MaHR) for them.
  • Our ablation study confirms the central role of local citation contexts in the learning process. It also shows the effectiveness of the Global training scheme over Base.
2 Related Work
BERT (Devlin et al., 2019) is an encoder-only pretraining model that adopts the Masked Language Modeling (MLM) objective. MLM masks tokens in a uniformly random fashion and predicts them, allowing the generation of learning signals bidirectionally. Some BERT variants were released to meet the requirements for masking a group of tokens. SpanBERT (Joshi et al., 2020) builds on this objective by masking random contiguous text spans. In the same direction, PMI-Masking (Levine et al., 2021) masks word n-grams based on their PMI (Pointwise Mutual Information) scores. Pretraining encoder decoders, e.g., BART (Lewis et al., 2020), combine the strengths of bidirectional learning of encoders with the autoregressive nature of decoders, capturing the local patterns of tokens within their generative capabilities.
The first citation-related task in natural language processing (NLP) has been citation impact prediction, where a paper’s future scientific impact is predicted on the basis of the number of times a paper gets cited after publication (Gehrke et al., 2003). Unlike the first approaches that relied on paper metadata and abstract, the recent work (van Dongen et al. (2020), Huang et al. (2022)) exploit the whole content of scientific papers to achieve the goal. Brody et al. (2006) aim to predict the future citations of a paper using web usage statistics. NNCP (Abrishami and Aliakbary, 2019) uses a SimpleRNN model to predict long-term citations using short-term citations. Bai et al. (2019) propose the Paper Potential Index based on a combination of manually acquired features. SChuBERT (van Dongen et al., 2020) leverages the entire contents of papers to accomplish the task. FGCCP (Huang et al., 2022) performs a fine-grained analysis to attribute citation frequencies to individual parts of papers.
Yu et al. (2012) learn citation relations through a meta path-based approach. Their approach combines authorship metadata with discriminative term features to calculate citation probabilities on the DBLP network. Tanner and Charniak (2015) combine LDA-Bayes with metadata features under a logistic regression classifier to recommend citations.
Similar to citation recommendation, the recent work of Luo et al. (2023) predicts provisions of the U.S. Code by pretraining RoBERTa (Liu et al., 2019) and LegalBERT (Chalkidis et al., 2020) on the curated dataset (PACER (Luo et al., 2023)) of the US federal court documents where each provision source text is given with its associated target citation. SciBERT (Beltagy et al., 2019) performs pretraining exclusively on scientific texts to learn global representations for scientific papers. SPECTER (Cohan et al., 2020) learns citation-aware global representations for scientific papers using a citation-based pretraining objective. SPECTER-produced representations introduced remarkable results in the paper classification and global citation recommendation tasks.
LCR has four benchmark datasets for evaluation. BERT-GCN (Jeong et al., 2020) introduced the FullTextPeerRead dataset, extended from the original PeerRead (Kang et al., 2018). Throughout this paper, we refer to the FullTextPeerRead dataset as PeerRead for brevity. An additional dataset is ACL-ARC (Bird et al., 2008), derived from the ACL Anthology Reference Corpus. We run our experiments on its ACL-200 subcategory, analogous to DualEnh (Medić and Snajder, 2020) and HAtten (Gu et al., 2022). Finally, Refseer (Huang et al., 2015) and ArXiv (Gu et al., 2022) are the largest benchmarks for this task.
BERT-GCN (Jeong et al., 2020) utilizes two encoders for citation recommendation. The first encoder generates local context embeddings using BERT, while the second one creates the graph embeddings of citation networks using a GCN model (Kipf and Welling, 2017). The approach combines these embeddings to produce representations for papers. It was evaluated exclusively on the PeerRead dataset.
DualEnh (Medić and Snajder, 2020) trains a Bi-LSTM model to leverage similarity between a target paper and its candidate papers. The target paper provides a context with a citation placeholder, and the model utilizes the titles and abstracts of candidate papers to calculate their semantic similarity scores. The authors calculate semantic and bibliographic scores to acquire the final recommendation scores as a weighted average. The bibliographic score is acquired by utilizing metadata such as author names and citation counts. The authors performed their experiments on the ACL-200 and Refseer datasets.
HAtten (Gu et al., 2022) uses a Hierarchical Attention Text Encoder and SciBERT-based Re-ranking scheme for LCR. It starts by pre-fetching potential candidate papers from a pool of citations. It accomplishes this filtering through a nearest neighborhood search between the local citation context plus the citing paper’s title and abstract (query text as a whole) and the title and abstracts from candidate target papers. In the re-ranking phase, the authors assign scores to candidate papers using a SciBERT model with a classification layer on top. HAtten achieves state-of-the-art results on all of the benchmark datasets.
Lastly, GM-s2orc-H (Buscaldi et al., 2024) proposes two approaches for predicting citation placeholders within a given context. Their first approach employs the GPT-2 model to determine whether a token could be part of a citation. The second approach performs a similar task using the BERT model, framing it as a Named Entity Recognition (NER) task. Their results confirm the superiority of the generative GPT-2 model over the second one. Although their results are not directly comparable to CiteBART due to differences in the task objectives, their findings highlight the advantages of generative models in citation-related tasks.
3 Methodology
We propose CiteBART, a novel pre-training strategy designed to predict citations within the contexts of scientific papers. We mask placeholder tokens, which replace ground-truth citations in the parenthetical author-date style, for the continual pre-training of a vanilla BART-base to generate the correct parenthetical author-date citation for a given context. CiteBART is trained on the benchmark datasets, learning to recommend citations during its generation process.
3.1 Custom BART Pre-training for LCR
BART Lewis et al. (2020) is a sequence-to-sequence model with an encoder and a decoder. It introduces a set of document corruption (denoising) schemes and then optimizes a reconstruction loss, the cross-entropy between the original document and the decoder’s outputs. The denoising transformations that are applied to the encoder during pre-training are as follows: Random token masking (similar to BERT), token deletion, text infilling (span masking with span lengths drawn from a Poisson distribution ($\lambda=3$)), sentence permutation, and document rotation with a randomly selected token leading the document.
We propose a citation learning strategy using BART. BART employs MLM similar to BERT. Additionally, to effectively reconstruct the masked contexts, it masks a span of $k$ tokens with a single mask. In return, it can predict multiple tokens for a single mask. Thus, CiteBART can generate complex parenthetical author-date citations after custom pre-training for citation tokens without requiring further architectural modifications.
We propose two training schemes for our approach: CiteBART-Base and CiteBART-Global. In CiteBART-Base, the model gets the masked context with the ground-truth citation as input. This setting tests the model’s performance in a local context-only situation (Table 1). With the underlying idea that good citation recommendation requires relating local citation contexts with the citing papers’ global information, such as titles and abstracts, we devised an innovative way to accomplish it. Inspiring from pre-training under the REALM framework (Guu et al., 2020), in CiteBART-Global, we append the citing paper’s title and abstract to the local context, allowing backpropagation through the global information that considers the pool of papers from the corpus. Specifically, we used the "$<$/s$>$" token designated by the pre-trained BART-base model as the separator.
Table 2: Statistics of LCR benchmarks.
Dataset Name ACL-200 PeerRead RefSeer Arxiv
Train Size 30,390 9,363 3,521,582 2,988,030
Validation Size 9,381 492 124,911 112,779
Test Size 9,585 6,184 126,593 104,401
# of Papers 19,776 4,837 624,957 1,661,201
Publication Years 2009-2015 2007-2017 -2014 1991-2020
3.2 Dataset Preprocessing
We conduct our experiments on the existing citation recommendation benchmarks of ACL-200, PeerRead, RefSeer, and Arxiv. Table 2 presents the statistics of these datasets. They provide citation contexts from various articles where all contexts have a target citation in the middle. The context sizes are in terms of characters, which causes some incomplete words at the start and end of the contexts.
The datasets originally include a "TARGETCIT" marker as a placeholder for citations within each context. We replaced these markers with "$<$mask$>$" tokens to align with our pretraining process. Additionally, to ensure CiteBART focuses solely on predicting target citations, we removed any non-target citations from all four datasets.
We encountered some issues during the preprocessing of ACL-200 and RefSeer. First, they include local contexts with author name conflicts in the citation tokens. For example, the "Petrović et al., 2010" citation token was incorrectly written as "Petrovic et al., 2010" in the target citation column of ACL-200. Another problem is the incorrect ordering of two-author citations. For instance, the local citation context provides the citation "Rivera and Zeinalian, 2016"; the paper metadata includes "Zeinalian and Rivera, 2016". There are also a few cases of incorrect citations. Moreover, there are some contexts with empty author names. We removed all these cases from the aforementioned datasets to ensure consistency.
After the preprocessing, we worked with the train and test sets. As CiteBART involves continual pre-training, we perform it on the training partition and evaluate the performance on the test partition. Table 3 shows the final statistics of our preprocessed datasets [^5] Please find information on token limits in Appendix A. including the training and test partition sizes for all the benchmarks.
Table 3: Statistics of the preprocessed datasets.
Dataset Name ACL-200 PeerRead RefSeer Arxiv
# of local contexts 63,365 16,669 3,739,189 3,205,210
Size of the training split 50,692 13,335 2,991,351 2,564,168
Size of the test split 12,673 3,334 747,838 641,042
# of removed contexts 403 0 39,577 0
# of unique citations 5,266 2,043 351,896 368,284
3.3 Metrics
To evaluate CiteBART, we used the Recall@$10$, Exact Match and Mean Reciprocal Rank metrics. The past works on citation recommendation have generally used Recall@$10$ and Mean Reciprocal Rank as evaluation metrics.
Recall@$10$ is the ratio of the correctly predicted items in the top k recommendations. The benchmark datasets have only one actual target for each context. Therefore, recall@$10$ measures whether the target citation matches any recommendations in top k.
Exact match (EM) calculates whether the first prediction of the model is the same as the target citation. It is the same as accuracy since there is only one ground-truth citation for each context.
Mean Reciprocal Rank (MRR) considers the position of the ground-truth label in a top-k ranked recommendation list. It is the mean of the reciprocal rank of the correctly recommended citation in the recommendation list. Thus, in Equation 1, $U$ corresponds to the total number of contexts in the dataset (test set size), and $i$ is the position of the ground-truth citation for context $u$ in the top-k results. We used $k$ as $10$ in our experiments.
$MRR=\frac{1}{U}\sum_{u=1}^{U}\frac{1}{rank_{i}}$ (1)
4 Experiments
We conducted our experiments on devices with NVIDIA RTX6000 Ada GPU and NVIDIA V100 GPU [^6] Please find information on training and evaluation times in Appendix B.. The following hyperparameters were utilized in all our experiments. The number of epochs was set to $15$, as the change in loss values between epochs became negligibly small beyond this point. Only the PeerRead Global dataset has been trained for 30 epochs since the generative model requires longer training for the relatively smaller PeerRead dataset. We employed a learning rate of $52e-$ 5 and an attention dropout rate of $0.12$. Given that BART is a generative model, we adjusted its generation parameters to produce outputs that align with our requirements. Specifically, we utilized the grouped beam search with $20$ beams and applied a diversity penalty of $1.5$ to generate more diverse results. The maximum number of generated tokens was $25$ since the generated citations should not exceed it. Apart from these specific modifications, we did not alter the architecture of the BART model.
4.1 Results
We report our results using Recall@10 (R@10), Exact Match (EM), and Mean Reciprocal Rank (MRR) and compare with the state-of-the-art approaches in Table 4. As can be seen from the table, CiteBART-Global outperforms others on the existing benchmarks except for the smallest PeerRead dataset, while the base scheme is still a strong baseline, surpassing BERT-GCN on PeerRead, DualEnh, and HAtten on Refseer. The table includes the best-reported results of HAtten with $2k$ pre-fetched candidates. As for DualEnh (Medić and Snajder, 2020), we chose their superior "DualEnh-ws" model for the comparison. BERT-GCN’s (Jeong et al., 2020) results are available only on the PeerRead dataset. None of the past works have provided their Exact Match scores.
As shown in Table 4, CiteBART-Global demonstrates its advantage over HAtten on Refseer most since Refseer includes more training contexts compared to ArXiv. Given that CiteBART is a generative model, access to a larger training set contributes to its improved results.
Table 4: Comparison with state-of-the-art on LCR benchmarks.
Model ACL-200 PeerRead Refseer Arxiv
R@10 EM MRR R@10 EM MRR R@10 EM MRR R@10 EM MRR
BERT-GCNa - - - 0.529 - 0.418 - - - - - -
DualEnh-ws 0.703 - 0.366 - - - 0.534 - 0.280 - - -
HAtten 0.633 - -c 0.757 - -c 0.454b - -c 0.439b - -c
CiteBART-Base 0.686 0.422 0.504 0.570 0.363 0.424 0.606 0.382 0.449 0.355 0.184 0.240
CiteBART-Global 0.739 0.417 0.513 0.669 0.430 0.502 0.652 0.404 0.479 0.502 0.230 0.305
  • a
    BERT-GCN performs evaluation by excluding the papers cited less than five times in each dataset.
  • b
    The reported results are based on a $10k$ subset of the test set.
  • c
    The authors did not report their MRR scores.
4.2 Qualitative Analysis
To provide insights into the working of CiteBART, we present some top $10$ prediction examples. We analyze four different scenarios shown in Table 5. Since CiteBART is a generative model, it is prone to hallucination. In the examples, the hallucinated predictions are designated with the * symbol. Furthermore, we provide another qualitative analysis on the performance of LLMs in LCR in Appendix C.
Table 5: Four example top-10 citation predictions using CiteBART. Due to space limitations, contexts and abstracts have been abbreviated. The hallucinated predictions are designated with the * symbol.
# Context Ground Truth Pretraining Dataset of the Model Dataset of the Example Predicted Citations
1 … Twitter. Previously, a series of NLP tasks have tried to utilize the social annotations like followers , emoticons and responses <mask> etc. two kinds of common social labels, i.e., hyper-links and hashtags are leveraged for … Hu et al., 2015 PeerRead Base PeerRead Base 1. Shang et al., 2015
2. Vinyals and Le, 2015
3. Baqapuri, 2015
4. Serban et al., 2015
5. Sordoni et al., 2015
6. Tan et al., 2015
7. Tan et al., 2014
8. Yin and Schutze, 2015 *
9. Dhingra et al., 2016
10. Tan et al., 2016
2 … Twitter. Previously, a series of NLP tasks have tried to utilize the social annotations like followers , emoticons and responses <mask> etc. two kinds of common social labels, i.e., hyper-links and hashtags are leveraged for … </s> TGSum: Build Tweet Guided Multi-Document Summarization Dataset </s> The development of summarization research has been significantly hampered by the … Hu et al., 2015 PeerRead Global PeerRead Global 1. Hu et al., 2015
2. Vinyals and Le, 2015
3. Bing et al., 2015
4. Tan et al., 2014
5. Dhingra et al., 2016
6. Xiao and Cho, 2016
7. Qu and Hovy, 2016 *
8. Bing et al., 2014 *
9. Lei et al., 2015
10. Qu and Zuidema, 2015 *
3 … in some latent space. There are many ways to structure G. The DCGAN <mask> uses fractionally-strided convolutions to upsample images instead of … </s> Gang of GANs: Generative Adversarial Networks with Maximum Margin Ranking </s> Traditional generative adversarial networks (GAN) and many of its variants are trained by minimizing the KL or JS-divergence loss … Radford et al., 2015 ACL-200 Global PeerRead Global 1. Kalchbrenner et al., 2014
2. Kalchbrenner and Blunsom, 2013
3. Sha and Pereira, 2003
4. Mikheev et al., 2013 *
5. Finkel et al., 2008
6. Mikheev et al., 1999
7. Gimpel and Smith, 2012
8. Kim et al., 2014
9. Blitzer et al., 2006
10. Henderson, 2004
4 … models to autoregressive models and stochastic variations of neural networks. Among them <mask> developed an approach for training a generative model with variational inference by performing … </s> Learning to Generate Chairs, Tables and Cars with Convolutional Networks </s> We train a generative convolutional neural network which is able to generate images of objects given object type, viewpoint … Rezende et al., 2014 Arxiv Global PeerRead Global 1. Rezende et al., 2014
2. Kusner and Hern’andez-lobato, 2016
3. Gregor et al., 2015
4. Mnih and Gregor, 2014
5. Doersch, 2016
6. Kusner and Hern’andez-lobato, 2015 *
7. Ioffe and Szegedy, 2015
8. Lamb et al., 2016
9. Salimans and Kingma, 2016
10. Salimans and Knowles, 2012
We first present an example context that is tested on a model pre-trained on the PeerRead Base dataset. It belongs to the test set of PeerRead Base and receives top $10$ citation predictions for the mask. As demonstrated below, the model fails to predict the correct citation in the top $10$ predictions. Actually, the ground-truth citation is the $18th$ entry in the ranked prediction list.
In a deeper analysis of the recommended citations for the first example, we bring up their connections with the ground-truth citation. The ground truth citation, "Hu et al., 2015", focuses on sentence-level semantics using convolutional neural networks (CNNs) with an application in dialogue generation. Similarly, the second prediction, "Vinyals and Le, 2015" leverages the sequential structure of sentences in dialogue systems. The fourth prediction, "Serban et al., 2015", also aims to model the hierarchical structure of sentences (utterances) for building an end-to-end dialogue system. The first prediction, "Shang et al., 2015," is still concerned with capturing sentence connections for a generative motivation. However, the primary reason for its top placement should be related to its experiments on Twitter data since the term Twitter appears in the local citation context. Analogously, the predictions $3,5,7$, and $9$ utilize Twitter as the data source. Lastly, the model may have proposed the entries $6$ and $10$ due to their overlaps in authors’ names with $7$.
The second example has the same context as the first one, but this time, the citing paper’s global information (title and abstract) is attached to it. Moreover, the model pre-trained on the PeerRead Global dataset makes the prediction, returning the ground truth citation in the first index. One can observe that the citations "Vinyals and Le, 2015", "Tan et al., 2015", and "Dhingra et al., 2016" still appear in the top-10 prediction list. There are also some hallucinated responses. The newly recommended "Bing et al., 2015" in the third position is also relevant since it tackles constructing sentences from fine-grained textual units.
The third example highlights our model’s cross-dataset generalization capability. We input a context from the PeerRead Global dataset into a model pre-trained on ACL-200 Global. The model fails to predict the correct citation as it is missing in the training dataset. Its predictions are NLP papers since ACL-200 is an NLP corpus. On the other hand, PeerRead includes both vision and text papers. The ground-truth citation, "Radford et al., 2015," focuses on image classification using CNNs, emphasizing unsupervised learning. Our analysis reveals that multiple predicted citations, among the top ten, are relevant to the ground-truth citation. For example, the papers in predictions 1 and 2 also employ CNNs but with a focus on sentence modeling. The papers from predictions 3 and 5 are about conditional random fields (CRFs). While their primary research areas differ significantly from the ground truth, terms such as ’conditional’ and ’random’ frequently appear in the ground truth paper. Moreover, the paper in Prediction 7 closely aligns with the ground-truth paper by strongly emphasizing unsupervised learning.
The fourth example emphasizes our model’s cross-dataset generalization capability from a different perspective. In this example, a model pre-trained on the Arxiv Global dataset manages to correctly predict the ground truth citation for a context from the PeerRead Global dataset. Upon closer inspection, we observed that this citation exists in both datasets but with different contexts. CiteBART-Global can predict the correct ground truth citation for an unseen context, leveraging another context citing the same reference.
4.3 Ablation Study
We conducted an ablation study to show different components’ contributions to the overall results. The analysis was carried out on the ACL-200 dataset. Table 6 shows the results for CiteBART with a model pre-trained on the ACL-200 Global dataset in $15$ epochs.
The first three experiments test the contribution of the local context, title, and abstract to the overall performance. First, we remove the local context to see the performance due to the global information-only training (#1 in Table 6). We discard the title and abstract in the second and third configurations (#2 and #3 in Table 6). The results show that excluding the local context brings about a sharp reduction in the performance metrics (a drop from $0.739$ to $0.588$ in Recall@10), confirming its decisive role in generating citations. On the other hand, removals of title or abstract do not lead to a statistically significant decrease in performance.
Table 6: Ablation study results on ACL-200 Global dataset under four different configurations.
Approach Training Input Recall@10 EM MRR
Base Context 0.686 0.422 0.504
Global Context + Citing Title & Abstract 0.739 0.417 0.513
1 No context Citing Title & Abstract 0.588 0.205 0.311
2 No title Context + Citing Abstract 0.731 0.415 0.509
3 No abstract Context + Citing Title 0.712 0.396 0.490
4 All-including Context + Citing Title & Abstract + Cited Title & Abstract 0.111 0.039 0.056
In the fourth ablation study, we further expand the global information with the cited paper’s title and abstract during pre-training (#4 in Table 6). The evaluation stays the same, feeding the local context with the citing paper’s title and abstract during inference. Contrary to expectations, adding the ground-truth paper’s global information during pre-training does not help; the model falls in its performance. This failure may be explained by the model learning to associate the citation token with the global information of both the citing and cited article in the training phase. However, lacking the cited paper’s global information in the test phase confuses the model’s predictions.
The previous studies (Medić and Snajder (2020), Gu et al. (2022)) utilize an all-including training and inference configuration where citing and cited paper’s global information is concatenated with the local citation context. Their pre-fetch and re-ranking pipeline is well-suited to this setup and benefits from it as the inference step also allows incorporating the cited paper’s title and abstract, which is not the case in a learning approach like ours’. Ideally, CiteBART-Global outperforms them without exploiting global information about the cited papers.
4.4 Taxonomy and Measurement of Hallucinated Citations
CiteBART, similar to other generative models, is prone to hallucination, occasionally producing citations that do not correspond to any real work. A generated citation is classified as hallucination if it is not present in the citation list of the dataset including the input context. Hallucinations in CiteBART are typically entity-error hallucinations or fabrications.
To measure the degree of hallucinations in LLM-generated responses, Li et al. (2024) propose two metrics, MaHR (macro hallucination rate) and MiHR (micro hallucination rate), respectively. While MaHR calculates the proportion of hallucinatory responses in all the responses (Equation 2), MiHR gives the average rate of hallucinations within each response (Equation 3).
$MaHR=\frac{Count(hallucinatory\ responses)}{n}$ (2)
$MiHR=\frac{1}{n}\sum_{i=1}^{n}\frac{Count(hallucinatory\ facts)}{Count(all\ facts\ in\ r_{i})}$ (3)
In LCR, MaHR represents the proportion of hallucinated citations across all generated citations. As the task is evaluated with top-$k$ predictions for each test instance, the total number of responses becomes $k*n$ where $n$ is the number of test instances. Thus, MaHR is the fraction of hallucinated citations among $k*n$ responses (Equation 4).
MiHR, on the other hand, measures the average hallucination rate in individual contexts. For example, each of $n$ contexts gets top-$k$ predictions and yields its hallucination rate, and MiHR is the average of these individual rates (Equation 5).
$MaHR=\frac{Count(hallucinated\ citations)}{k*n}$ (4)
$MiHR=\frac{1}{n}\sum_{i=1}^{n}\frac{Count(hallucinated\ citations\ in\ context_{i})}{k}$ (5)
In LCR, as each context gets top-$k$ predictions, the number of facts in each response is fixed with $k$ (the denominator in Equation 5), which makes MaHR and MiHR produce identical results.
In addition to MaHR (or MiHR), we propose the following metrics to pinpoint hallucination behavior. Each metric targets a type of hallucination we categorized by examining hallucinations versus ground truth citations for given contexts.
  • Incorrect year (all-names-GT): The generated citation fully matches the author(s) in the ground truth citation while failing to match the publication year.
  • Partially correct author list (one-name-GT): One of the two author names is correct, and the generated year may or may not be correct in these cases.
  • Correct year with incorrect authors (year-GT): Some hallucinations match the year of the ground truth citation, even if the author names are incorrect.
  • wrong-format: If the generated citation’s format does not conform to the parenthetical author-date citation style, it is considered a wrong-format hallucination. These types of hallucinations happen very rarely.
  • other-hal: The other types of hallucinations that do not belong to any of the above types belong to this category. There is no overlap with any part of GT in these hallucinations.
Additionally, we term the aggregation of the hallucinations corresponding to partially correct responses MaHR-partial and calculate it using Equation 6. Lastly, we relate MaHR with MaHR-partial using Equation 7.
$MaHR\mhyphen partial=all\mhyphen names\mhyphen GT+one\mhyphen name\mhyphen GT+year\mhyphen GT$ (6)
$MaHR=MaHR\mhyphen partial+wrong\mhyphen format+other\mhyphen hal$ (7)
Table 7 presents the results of the hallucination metrics for the CiteBART-Global models. To observe the effect of the $k$ value, we performed each analysis with top-3, top-5, and top-10 generated predictions, respectively. The results conclude that MaHR-partial accounts for almost half of the hallucinations in the top 3 predictions, which implies that when the model is forced to make fewer predictions, its hallucinations do not deviate much from the ground truth. The proportion gradually diminishes in the top-5 and top-10 predictions. Interestingly, on Refseer and Arxiv Global, the incorrect year (all-names-GT) hallucination, which is the closest to the ground truth, decreases with increasing $k$ values. In overall performance, the ACL-200 Global dataset gives the lowest hallucination rates all over the $k$ values. Arxiv Global is the second best, with very close scores to ACL-200 Global.
Table 7: Results for proposed hallucination metrics on Global datasets for top-3, top-5, and top-10 predictions. Metric values are shown as percentages (%). The best values are shown with bold.
Metrics ACL-200 PeerRead Refseer Arxiv
Top-3 Top-5 Top-10 Top-3 Top-5 Top-10 Top-3 Top-5 Top-10 Top-3 Top-5 Top-10
all-names-GT 0.63 0.54 0.68 0.63 0.59 0.85 1.21 1.07 1.01 1.01 0.85 0.80
one-name-GT 0.24 0.29 0.44 1.31 1.06 1.21 0.63 0.75 0.82 0.43 0.56 0.65
year-GT 1.03 1.56 2.48 1.72 3.08 5.95 0.50 0.84 1.48 0.55 0.99 1.94
MaHR-partiala 1.89 2.39 3.60 3.66 4.73 8.01 2.34 2.66 3.31 1.99 2.40 3.39
wrong-format 0.02 0.02 0.08 0.00 0.07 0.28 2.18e-5 4.97e-5 0.01 1.35e-5 2.12e-5 0.01
other-hal 2.20 4.00 9.02 4.36 7.26 15.02 2.94 5.25 10.05 2.66 4.74 9.95
MaHRb 4.12 6.42 12.69 8.02 12.06 23.31 5.28 7.91 13.37 4.64 7.14 13.35
Table 8 reports the values of some extended metrics built upon MaHR:
  • top-k-match-MaHR: This metric considers hallucinated predictions only when one of the other predictions in the same top-k group matches the ground truth (GT).
  • exact-match-MaHR: This metric is similar to top-k-match-MaHR but specifically focuses on the cases where the exact match occurs (the first prediction is correct).
These metrics approach the problem differently by examining the hallucination tendency when the model can hit the ground truth citation in its top-k predictions. In other words, the research question is whether the model suffers less from the hallucination given the correct prediction in the top-k list (when the model knows the answer). The results confirm this hypothesis as top-k-match-MaHR and exact-match-MaHR are different from MaHR in a statistically significant way with $p<0.001$. Furthermore, Arxiv Global is the best model to mitigate hallucinations when it hits the ground truth, outperforming others in the hallucination rates.
Table 8: Results for extended MaHR metrics on Global datasets for top-3, top-5, and top-10 predictions. Metric values are shown as percentages (%). The best values are shown with bold.
Metrics ACL-200 PeerRead Refseer Arxiv
Top-3 Top-5 Top-10 Top-3 Top-5 Top-10 Top-3 Top-5 Top-10 Top-3 Top-5 Top-10
MaHR 4.12 6.42 12.69 8.02 12.06 23.31 5.28 7.91 13.37 4.64 7.14 13.35
top-k-match-MaHR 2.54 4.55 9.69 5.05 8.15 16.53 2.76 4.74 8.60 1.79 3.12 6.28
exact-match-MaHR 2.40 3.81 7.08 4.76 6.94 12.42 2.46 3.97 6.58 1.48 2.33 3.95
4.5 Qualitative Analysis on Hallucinations
In this section, we provide additional examples to illustrate the types of hallucinations (Table 9). The first example shows an ideal scenario with no hallucinations in the top-10 prediction list. The other examples, except the last, depict different types of hallucinations. The last example showcases the cross-dataset generalization capability of CiteBART. Due to space limitations, contexts and abstracts have been abbreviated. Hallucinated predictions are designated with the * symbol.
Table 9: Examples of hallucination categories. (a) No hallucination in any of the top-10 predictions. (b) Hallucinated publication years in the fourth, sixth, and ninth predictions. (c) Hallucinated author name in the sixth prediction. Fabricated author list in the ninth prediction. (d) Hallucinated author name in the fifth prediction. (A typo in the first author’s name). (e) Hallucinated author name in the sixth prediction (A single letter as the first author name). (f) In this cross-dataset example, CiteBART predicts a citation that has the same author name as the ground truth while in a different citation format and publication year.
Context Ground Truth Pretraining Dataset of the Model Dataset of the Example Predicted Citations
(a) … exploits similarity on the target side in another language by extracting source phrases that share common translations <mask> , but recent approaches have combined this approach with source phrase … </s> Example-based Paraphrasing for Improved Phrase-Based Statistical Machine Translation </s> In this article, an original view on how to improve phrase translation estimates is proposed. This proposal is … Bannard and Callison-Burch, 2005 ACL-200 Global ACL-200 Global 1. Callison-Burch et al., 2006
2. Koehn et al., 2003
3. Irvine and Callison-Burch, 2014
4. Bannard and Callison-Burch, 2005
5. Quirk et al., 2004
6. Mirkin et al., 2009
7. Irvine and Callison-Burch, 2013
8. Koehn and Knight, 2002
9. Schroeder et al., 2009
10. Koehn and Knight, 2003
(b) … supertags, the supertagger re-analyses the sentence with a more relaxed beam (adaptive supertagging). A* Parsing <mask> a) introduce A* parsing for PCFGs. The parser maintains a chart and an agenda, which is a priority queue of … </s> A* CCG Parsing with a Supertag-factored Model </s> We introduce a new CCG parsing model which is factored on lexical category assignments. Parsing is then simply a deterministic … Klein and Manning, 2003 ACL-200 Global ACL-200 Global 1. Klein and Manning, 2003
2. Auli and Lopez, 2011
3. Ait-Mokhtar and Chanod, 1997
4. Ait-Mokhtar and Chanod, 2005 *
5. Pauls et al., 2009
6. Pauls et al., 2006 *
7. Ait-Mokhtar and Chanod, 2006 *
8. Och, 2003
9. Aitouni et al., 2006 *
10. Clark and Curran, 2004
(c) … Google Analogy Test Set, which contains 14 types of relations with a varying number of instances per relation <mask>, the gger Analogy Test Set , which contains 40 relations with 50 instances per relation, and the ffVec Test Set … </s> Probabilistic Relation Induction in Vector Space Embeddings </s> Word embeddings have been found to capture a surprisingly rich amount of syntactic and semantic knowledge. However, it is not … Mikolov et al., 2013 PeerRead Global PeerRead Global 1. Vylomova et al., 2015
2. Valenzuela-escárcega et al., 2015
3. Abadi et al., 2016
4. Heinsohn, 2013
5. Holzmann and Risse, 2017
6. Valenzuela-escárárcega et al., 2015 *
7. Davies et al., 2015 *
8. Dinu et al., 2014
9. Holzmann and Riedl, 2016 *
10. Gaunt et al., 2016
(d) … produces a false positive rate of 0.0027, as noted above, but in a situation where 3 key-value items were being stored per n-gram on average, this error rate would in fact require a storage cost of 60 bits per original n-gram. 2.2.2 Bloomier Filters More recently, <mask> have proposed an approach to storing large language models which is based on the Bloomier Filter technique of OTHERCIT. Bloomier Filters generalize the Bloom Filter to allow values … Talbot and Brants, 2008 ACL-200 Base ACL-200 Base 1. Talbot and Brants, 2008
2. Talbot and Osborne, 2007
3. Lavoie and Rambow, 1997
4. Pennacchiotti and Pantel, 2009
5. MTalbot and Brants, 2008 *
6. Galanis and Androutsopoulos, 2010
7. Lavoie and Rambow, 2009 *
8. Pennacchiotti and Pantel, 2006
9. Mintz et al., 2009
10. Talbot et al., 2011
(e) … signature generators can be mislead into generating bad signatures; specifically higher false negative rates. Shield <mask> , Vigilante, DACODA, and our own work, all attempt to work around such problems by directly deriving … </s> A lightweight end-to-end system for defending against fast worms </s> The vulnerabilities which plague computers cause endless grief to users. Slammer compromised millions of hosts in minutes; a hit-list worm … Wang et al., 2004 Refseer Global Refseer Global 1. Wang et al., 2004
2. Cui et al., 2007
3. Brumley et al., 2006
4. Brumley et al., 2004 *
5. Dasgupta et al., 2004
6. W et al., 2004 *
7. Shavitt and Tankel, 2003
8. Shavitt and Tanenbaum, 2005 *
9. Daswani and S, 2007 *
10. Chen and Wagner, 2007
(f) … tab while waiting for the original one to load, i.e., tab switching. More recently, a Web navigation study by <mask> found their participants using multiple windows frequently, enabling them to compare search results … </s> Parallel Browsing Behavior on the Web </s> Parallel browsing describes a behavior where users visit Web pages in multiple concurrent threads. Web browsers explicitly support this by providing tabs. Although parallel browsing … Weinreich, 2006 ACL-200 Global Refseer Global 1. Weinreich et al., 2008
2. Nakagawa and Uchimoto, 2007 *
3. Weinreich et al., 2010 *
4. Navigli and Crisafulli, 2010
5. Nakashole et al., 2012 *
6. Webber et al., 2003 *
7. Lin and Bilmes, 2011
8. Resnik and Smith, 2003
9. Stoica and Hearst, 2004 *
10. Lin and Bilmes, 2008 *
5 Discussion and Conclusion
CiteBART is distinctive as it performs LCR by end-to-end learning. On the other hand, the recent approaches adopt pre-fetch and re-rank pipelines where their system first retrieves a set of papers and then ranks the retrieved by matching queries (citing papers’ titles and abstracts, local citation contexts) with candidate papers’ representations (cited papers’ titles and abstracts). While our model does not use global information about cited papers during testing, these systems require titles and abstracts of the cited papers for inference. In CiteBART-Base, we rely solely on local citation contexts, while CiteBART-Global incorporates the citing paper’s global information to make predictions. CiteBART-Global achieves state-of-the-art performance on LCR benchmarks except for the FullTextPeerRead dataset, which is quite small to see the advantage of generative pre-training.
CiteBART can still be fine-tuned for any downstream task. We hypothesize that it should perform better in downstream tasks involving citations and scientific papers than other language models without citation-specific learning signals during pre-training, an area we intend to explore in future work. Furthermore, with the release of new citation recommendation datasets, it will be sufficient to continually pre-train the model to acquire knowledge about the new scientific papers with no need to pre-train from scratch.
We comment on the advantages of using BART over encoder-based pre-training models such as RoBERTa. BART’s MLM objective is flexible and allows the masking of all the tokens in the parenthetical author-date style. RoBERTa cannot add citation tokens to its vocabulary by its MLM. Moreover, in the prediction phase, constraining predictions to citation tokens for RoBERTa is not straightforward. While BART is prone to hallucination, its capabilities significantly enhance LCR performance.
Furthermore, our comprehensive hallucination analysis sheds light on the hallucination behavior, MaHR-partial taking up significant proportions (almost half of the hallucinations in the top 3 predictions), which implies that all the hallucinations should not be rejected beforehand but show signs of promising zero-shot capabilities as MaHR-partial is the aggregation of partially correct hallucinations that are correct in all the author names, single author names, and year, respectively. The hallucinations that are (partially) correct in the author names may be useful for finding suggested reading material along with the ground truth paper as they reveal relevant authors. Another finding is that when the prediction is successful in the top-k list, the hallucination tendency in the other predictions drops significantly, the Arxiv Global trained model being the most advantageous, highlighting that the largest model also shows good traits in mitigating hallucinations. The evidence on hallucinations in this study may also lead to hallucination analyses in other domains that clear up generative models’ hallucination landscape.
As shown in our ablation study, extending the local citation context with both the citing and cited paper’s title and abstract during the continual pre-training does not produce a better result, which can be evaluated counter-intuitive as one has all the information to learn a citation relationship. The missing global information for the cited paper in the test phase complicates finding out the associated citation token.
For future work, we plan to investigate further the all-including configuration given in the ablation study. Conceptually, exploiting the cited paper’s title and abstract during the continual pre-training should have been complementary. However, the empirical evidence proves the contrary. More sophisticated masking strategies besides citation token masking should connect the dots by combining the information from the citing paper’s title and abstract, local citation context, and the cited paper’s title and abstract. Additionally, we should investigate the potential solutions to the citation-specific hallucinations and tackle a way to reduce the number of hallucinated recommendations in the top k.
Ethics Statement
CiteBART is a tool to support the scientific community in paper writing; it in no way replaces a researcher or alternates the thoughtful process of choosing the most appropriate references to cite in a local context.
Acknowledgments
The Scientific and Technological Research Council of Türkiye (TUBITAK) supported this research with the 2219 fellowship awarded to the second author as a visiting scholar at the University of Edinburgh School of Informatics.
We primarily used the hardware purchased by the project supported by the Council of Higher Education (YÖK) under ADEP grant number 2022IYTE-3-0027 for our experiments. They were partially run at TUBITAK ULAKBIM, High Performance and Grid Computing Center (TRUBA resources).
Appendix A Token Limits
Before pre-training with citation objectives, we ensured that each context has its "$<$mask$>$" token in its middle position after tokenization. Another critical aspect was the determination of correct lengths for citation contexts. We limited citation contexts in each dataset to an optimal number of tokens to avoid increasing time and memory costs. An exploratory analysis of context lengths shows that the contexts of ACL-200 and Peerread are significantly longer than those of the other datasets. After tokenization, we observed that $200-400$ tokens were optimal for all base datasets. This limit allows sufficiently long contexts without a need for excessive amounts of padding tokens. As an exception, ACL-200 has $607$ contexts that exceed the $400$ limit. We have shortened them to the $400$ token limit as they correspond to a small proportion of the whole number of contexts and also because the number of discarded tokens is negligible.
Table 10: Maximum token limits for the preprocessed datasets.
Dataset Name Base Token Limit Global Token Limit
ACL-200 400 350
FullTextPeerRead 400 350
Refseer 200 350
Arxiv 300 350
For each global dataset, we chose the token limit as $350$. Since abstracts require a higher number of tokens, we limited the local context sizes to $100$ for the global versions of the datasets. We also ensured that there are $50$ tokens each on the left and right sides of the <mask> tokens. We used a token limit of $200$ for abstracts for all datasets since most abstracts can fit into it. Table 10 shows the maximum token limits for both the base and global training schemes.
Appendix B Training and Evaluation Times
We conducted our experiments on devices with NVIDIA RTX6000 Ada GPU and NVIDIA V100 GPU for Global and Base datasets, respectively. For global datasets, the pre-training for Peerread and ACL-200 lasts for $2$ and $6$ hours, respectively. The larger datasets, Arxiv and Refseer, take up to $8-9$ days since they have similar sizes. For base datasets, the training for the smaller datasets, Peerread and ACL-200, lasts for 8 and 20 hours, respectively. The larger datasets, Arxiv and Refseer, take up to 14-15 days. However, we believe these relatively longer times are the result of training on the device with NVIDIA V100 GPU.
Our evaluation of the corresponding test sets takes considerable time since generating the top 10 predictions for each example is resource-intensive. Especially with our limited hardware resources, acquiring the results on the larger datasets takes up to 2 days. The smaller datasets require less time, 20 minutes for Peerread and 2 hours for ACL-200. We performed our evaluations on the device with NVIDIA RTX6000 Ada GPU.
The issue of slow evaluation for larger datasets is not exclusive to our work. Gu et al. [2022] reported their results using only a smaller subsection ($10K$) of the test sets due to long evaluation times.
Appendix C Qualitative Analysis on Large Language Models’ Performances in LCR
We conducted experiments on a Large Language Model (LLM) to evaluate its performance in local citation recommendation. We prompted the open-source "Llama-2-70b-chat" model for our trials. In each prompt, we first list a set of citation tokens ($200$, due to the limits of chat windows) from our dataset, followed by a few examples of masked contexts with the corresponding ground truth mask values. Subsequently, we ask the model to fill in the mask for a new context by selecting a citation from the initially provided list.
Figure 1: Prompt examples on a Large Language Model for Base dataset.
We present four examples in Figures 1 and 2 to illustrate the workings of the base and global pre-training schemes, respectively. Due to space constraints, we partially display the list of citations, example contexts, and citing abstracts in the prompts. Each example consists of three parts: the prompt, the LLM’s answer, and the ground truth value of the masked citation token provided at the end of the prompt.
Figure 2: Prompt examples on a Large Language Model for Global dataset.
Figure 1 includes a correct prediction in Part (a) and an incorrect one in (b). Indeed, the correct prediction is the only successful example in several trials using the base approach. The model responds to the prompt by "Shwartz et al., 2016" explaining its choice. On the other hand, the model fills in the mask by "Bahdanau et al., 2016" in Part (b), where "Bluche, 2016" is expected. Its reasoning sheds light on its wrong choice as it strongly associates the term "attention-based mechanisms" in the local context with Bahdanau et al.’s seminal paper on attention-based sequence modeling.
In Figure 2, Part (a) presents a successful example based on the global dataset where the prompt includes the citing paper’s title and abstract with the local citation context. The LLM generates the correct citation without an explanation, unlike other predictions. The second example in Part (b) belongs to an incorrect prediction, yet the LLM makes a plausible choice here, judging from its grounding. We can conclude from the observed behavior that LLMs need custom pre-training for the citation tokens to perform well in the task of local citation recommendation.
Our further trials with LLMs demonstrate that they tend not to restrict their predictions to the provided list of citations but to recommend the best choice based on their prior knowledge. They also exhibit a known deficiency. They sometimes ask for confirmation when they provide an answer, and even if you confirm, they lean towards changing the answer. In conclusion, they suffer from hallucinations.